# Qubit Mapping and Routing tailored to Advanced Quantum ISAs: Not as Costly as You Think

### v1

@@ Abstract.
To address this, we propose Canopus, a unified qubit mapping/routing framework applicable to diverse quantum ISAs. Built upon the canonical representation of two-qubit gates, Canopus centers on qubit routing to perform deep co-optimization in an ISA-aware approach. Canopus leverages the two-qubit canonical representation and the monodromy polytope to model the synthesis cost for more intelligent $\mathrm{SWAP}$ insertion during the routing stage. We also formalize the commutation relations between two-qubit gates through the canonical form, providing a generalized approach to commutativity-based optimizations. Experiments show that Canopus consistently reduces routing overhead by 15%-35% compared to state-of-the-art methods across different ISAs and topologies. Our work also presents a coherent method for co-exploration of program patterns, quantum ISAs, and hardware topologies. We have for the first time demonstrated that advanced quantum ISAs can be efficiently utilized within a unified routing framework, paving the way for more effective co-design of quantum software and hardware.
@@ 1. Introduction
Holistic benchmarks of quantum computers such as quantum volume (cross2019validating) are predicated on concurrent advancements in both hardware and software. Recently, numerous systematic techniques regarding compiler optimization and architectural support have been presented to push the limit of hardware performance. Quantum compilers play a pivotal role in this process, translating high-level programs into executable single-qubit (1Q) and two-qubit (2Q) gates on realistic quantum hardware. This typically involves several stages: (1) compiling programs into basic quantum gates, (2) performing hardware-agnostic (logical-level) circuit optimization, (3) resolving backend topology constraints via qubit placement and routing, and (4) converting circuits to native gates for final optimization and scheduling. The primary goal of compiler optimization is to lower the 2Q gate count and circuit depth while resolving backend constraints, with a particular emphasis on 2Q gates due to their significantly higher error rates compared to 1Q gates.
@@
Most studies on qubit routing rely on a simplified routing model, where circuit cost is quantified by the $\mathrm{CX}$-based gate count and circuit depth while each $\mathrm{SWAP}$ gate is unrolled into three $\mathrm{CX}$ gates according to the textbook pattern $\mathrm{SWAP}_{q_{0},q_{1}}=\mathrm{CX}_{q_{0},q_{1}}\mathrm{CX}_{q_{1},q_{0}}\mathrm{CX}_{q_{0},q_{1}}$. However, this $\mathrm{CX}$-centric view is misaligned with the physical reality of modern quantum devices. Although quantum algorithms are typically expressed in terms of $\mathrm{CX}$ gates, the underlying hardware may not execute native $\mathrm{CX}$-equivalent gates, nor does this gate/circuit cost quantification method accurately reflect the true operational cost. Indeed, beyond the native support for $\mathrm{CX}$-equivalent gates (e.g., $\mathrm{CZ}$ (krantz2019quantum), Cross-Resonance (rigetti2010fully), Mølmer-Sørensen (bruzewicz2019trapped)), modern quantum hardwares increasingly feature diverse native 2Q basis gates in recent years. These alternative basis gates, or the abstracted instruction set architectures (ISAs) can be more powerful than $\mathrm{CX}$-equivalent gates in terms of synthesis capabilities and fidelity of realization, such as $\mathrm{\sqrt{iSWAP}}$ (huang2023quantum), the $\mathrm{iSWAP}$-family and $\mathrm{CX}$-family gates (mckinney2024mirage; qiskitXXDecomposer), and heterogeneous/combinatorial basis gates (peterson2022optimal; mckinney2024mirage). With such ISAs, $\mathrm{SWAP}$ can be implemented with a lower cost than three $\mathrm{CX}$ gates or even be natively realized with high fidelity (wei2024native; chen2025efficient; nguyen2024programmable). Therefore, the simplified routing model completely ignores the backend ISA properties, severely limiting the potential of compiler optimization. Furthermore, the absence of systematic compiler optimization methods across these diverse (even complex, heterogeneous) ISAs has prevented the community from fully exploiting their power and exploring the rich software-hardware co-design space.
@@
The main ideas of Canopus are as follows: ① Considering the subsequent synthesis overhead with native basis gates after the qubit routing stage could unlock significant optimization opportunities. For instance, a $\mathrm{SWAP}$ inserted after a 2Q block with the same qubit pair acted on could lead to a better synthesis result when rebased to the backend ISA. ② Expanding the quantum ISA is crucial to boost the performance of real-world quantum applications. For example, the fractional $\mathrm{ZZ}(\theta)$ gate set widely adopted by hardware vendors (e.g., IBM (ibmFractionalGates), Quantinuum (quantinuumArbitraryAngleGates), IonQ (ionqPartialGates)) enables more efficient execution of chemistry simulation kernels within which many 2-local Pauli rotations are involved. The combination of $\mathrm{CX}$ and $\mathrm{iSWAP}$ gates have been demonstrated to benefit stabilizer circuits to protect error-corrected qubit information (zhou2024halma). ③ The monodromy polytope theory (peterson2020fixed) based on the canonical representation of 2Q gates (zhang2003geometric) provides a formal, universal, and quantitative description of the 2Q synthesis cost for generic quantum ISAs. This formalism enables unified compiler optimization. With these insights in mind, Canopus performs intelligent $\mathrm{SWAP}$ insertion during qubit routing to holistically minimize post-mapping circuit cost (in terms of both gate count and depth) given any quantum ISA, thus performing deep routing-synthesis co-optimization and resulting in significantly lower routing overheads. Importantly, although Canopus is ISA-aware, it always operates on the canonical-form circuits, and the gate/circuit cost quantification via monodromy polytopes is independent of the specific ISA rebase implementation. Thus Canopus offers an LLVM-style optimization.
Experimental results demonstrate that Canopus consistently provides 15%-35% reduction (in terms of both gate count and depth) of routing overhead compared to other SOTA methods across representative quantum ISAs, including the conventional CX ISA. This cross-ISA comparison also reveals some consistent or program-specific or topology-specific guidelines for hardware-software co-design. Source code and data are available via the Github link. Our work makes the following key contributions:
❶We utilize the canonical 2Q gate representation and the monodrome polytope to quantify costs of 2Q gates and the overall circuit. This formal approach accurately guides synthesis-routing co-optimization and cross-ISA evaluation.
@@
❺Our case studies, including the real-machine QFT kernel execution and the end-to-end QEC circuit simulation, unequivocally showcase Canopus’ superiority in both near-term and fault-tolerant applications. For example, on the task of mapping QFT on 1D chain topology, Canopus finds the provably optimal routing scheme, surpassing the results previously reported as optimal in prior work (zhang2021time); and experiments on IBM’s QPUs suggest that Canopus leads to an average 2.1x fidelity improvement compared to Qiskit.
@@ 2.2. Canonical description of 2Q gates
Figure 3. Geometric illustration of canonical gates confined to the Weyl chamber. For visualization convenience, herein the Weyl chamber is confined to $\left\{\frac{\pi}{4}\geq x\geq y\geq z\geq 0\right\}\cup\left\{\frac{\pi}{4}\geq\frac{\pi}{2}-x\geq y\geq z\geq 0\right\}$, equivalent to the canonical coefficient convention $\left\{(a,b,c)\,|\,\frac{1}{2}\geq a\geq b\geq|c|\right\}$ throughout this paper.
@@
(1) $\mathrm{Can}(a,b,c):=e^{-i\frac{\pi}{2}(a\,XX+b\,YY+c\,ZZ)},\,\frac{1}{2}\geq a\geq b\geq\lvert c\rvert$
sandwiched by local 1Q gates such that we say $U$ is locally equivalent to the canonical form $\mathrm{Can}(a,b,c)$.
@@ 2.3. Gate realization cost on hardware
The transformed circuits via qubit routing will be ultimately converted into basis gates for execution on hardware. Basis gates refer to those natively implemented and calibrated on physical platforms. Typical native gates in superconducting are $\mathrm{CR}$ (rigetti2010fully), $\mathrm{CZ}$, and $\mathrm{iSWAP}$ gates (krantz2019quantum). The realization cost of basis gates involves multiple aspects, including the benchmarked fidelity, gate duration, calibration efficiency, etc. For example, gates with shorter duration are more likely to achieve high fidelity, as qubit decoherence dominates the noise source; although some gate schemes can now implement more basis gates (chen2025efficient; nguyen2024programmable), those with simpler pulse control are more likely to be calibrated with high precision, such as the $\mathrm{iSWAP}$ family on flux-tunable transmons.
2Q gates that are not natively implemented must be synthesized by native gates. Their realization cost is determined by what basis gates are required for synthesis. For example, any 2Q gate can be minimally synthesized by 3 $\mathrm{CX}$ gates, except for $\mathrm{Can}(a,b,0)$ for which the required $\mathrm{CX}$ count is 2. Conventionally, $\mathrm{SWAP}$ is regarded as 3 times that of $\mathrm{CX}$ realization cost, while it can also be synthesized by “1 $\mathrm{CX}$ + 1 $\mathrm{iSWAP}$” or “3 $\mathrm{\sqrt{iSWAP}}$” gates. The monodromy polytope theory was proposed to determine the optimal synthesis cost for any 2Q gate given a specific set of basis gates through analysis of local invariants of canonical gates (peterson2020fixed). By this method, the set of gates realizable by a specified number of 2Q gates from the basis set, with arbitrary 1Q gates, corresponds to a polytope within the Weyl chamber. For instance, the polytope for 2 $\mathrm{\sqrt{iSWAP}}$ gates is a tetrahedron confined to $\left\{1/2\geq a\geq b+\lvert c\rvert\right\}$ (huang2023quantum).
@@ 3. Motivation
Limitations of conventional qubit routing models.
@@
Co-optimization as the key to unlocking the superiority of advanced ISAs.
@@
Prior efforts have been narrowly focused on local 2Q or multi-qubit synthesis tasks (huang2023quantum; tang2024quantum) or brute-force numerical optimizations (davis2019heuristics; younis2021qfast). Such rebase passes are tailored to a specific quantum ISA and fail to provide clear benefits when applied to advanced ISAs in realistic workloads (kalloor2024quantum). This has led to a critical question lingering in the community: “Are these more expressive, noise-resilient ISAs actually better?” Recently there have been attempts to harness the properties of advanced ISAs, although through manual and unsophisticated design, such as the $\mathrm{\sqrt{iSWAP}}$-based routing-synthesis optimization (mckinney2024mirage) and the $\mathrm{CX}$-$\mathrm{iSWAP}$ based routing for defect mitigation (zhou2024halma). In our work, we highlight that collaborative compiler optimization, especially at the stage following logical-level circuit optimization and followed by the final ISA rebase pass, is a key to fully exploit capabilities of those powerful ISAs: First, high-level algorithms are expressed in the $\mathrm{CX}$ representation, which then undergo template-based and peephole optimizations that are highly sophisticated and tailored for $\mathrm{CX}$-based circuit patterns (e.g., commutativity, Clifford equivalence); Second, the gap between the naïve qubit routing model and backend ISA properties apparently leaves a large untapped co-optimization space for both qubit routing and ISA rebase. Thus we are committed to validating this point through a systematic ISA-aware routing framework.
The “Tower of Babel dilemma” for utilizing diverse ISAs.
The proliferation of diverse quantum ISAs—from monolithic to complex, heterogeneous basis gate sets supported by various physical platforms—has created a “Tower of Babel dilemma” in the architecture and systems community. Developing bespoke compiler optimizations for each unique hardware backend is unsustainable, leading to the same software fragmentation that we have encountered in classical computing. Consequently, it is important to seek a unified approach that can effectively handle various platform-specific abstractions resembling the LLVM compiler (Lattner2004). The recently proposed monodromy polytope theory (peterson2020fixed), for example, provides the the method for optimal analysis of ISA synthesis capabilities. Specifically regarding the circuit-level compiler optimization, the monodromy polytope with canonical 2Q gate representations offers a unified approach to evaluating circuit cost and modeling routing-synthesis co-optimization. Building on this, Canopus offers an elegant and unified solution to the Tower of Babel dilemma at the compiler level.
Coherent cross-ISA, topology, and program pattern co-exploration.
@@ 4.1. Overview
The overall routing procedure of Canopus is illustrated in Figure 4. The input circuit is rebased to $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ before being fed to the routing pass. All processes operate on the directed acyclic graph (DAG) representation of the circuit. Canopus integrates the ISA-specific synthesis cost model into its $\mathrm{SWAP}$ search process and determines the best $\mathrm{SWAP}$ at each route step. The routing cost is efficiently computed by a formal analysis of 2Q canonical forms, without the ISA rebase step. Thus, the output is still a circuit DAG represented in $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ with inserted $\mathrm{SWAP}$ gates.
Notably, Canopus inherits the basic concepts and data structures introduced in Sabre (li2019tackling), which is one of the industrial-standard qubit mapping algorithms. Given the input circuit DAG, Canopus first attempts to map 2Q gates layer by layer via extracting the front layer $F$, peeling executable gates and searching $\mathrm{SWAP}$ gates to minimize the routing cost according to a heuristic cost function. On the backbone of Sabre, we further introduce several key data structures such as the last mapped layer $L$ and the wire duration records $D$, to support the efficient implementation of ISA-aware routing and reducing both gate count and depth related routing overhead.
@@
As introduced in Section 2.3, given any basis gate set, the synthesis cost of a target 2Q gate can be exactly computed through the monodromy polytopes (peterson2020fixed). This cost (which basis gates are sufficient for synthesis) only depends on the canonical coefficients of the target gate. For example, Figure 5 illustrates various polytopes for the gate set $\bigl\{\mathrm{\sqrt{iSWAP}},\,\mathrm{ECP}\bigr\}$. In practice, the costs of each basis gate are pre-defined, thus the whole polytope helps decide the optimal synthesis scheme we should prioritize with the minimal circuit cost. For example, if $\mathrm{\sqrt{iSWAP}}$ and $\mathrm{ECP}$ have the same unit cost, the $\mathrm{SWAP}\sim\mathrm{Can}\bigl(\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)$ gate realization will prioritize the “1 $\mathrm{\sqrt{iSWAP}}$ + 1 $\mathrm{ECP}$” combination; if $\mathrm{ECP}$ cost is set to be more than twice that of $\mathrm{\sqrt{iSWAP}}$, the $\mathrm{SWAP}$ realization prioritizes the “3 $\mathrm{\sqrt{iSWAP}}$” pattern.

### v3

@@ Abstract.
To address this, we propose Canopus, a unified qubit mapping/routing framework applicable to diverse quantum ISAs. Built upon the canonical representation of two-qubit gates, Canopus centers on qubit routing to perform deep co-optimization in an ISA-aware approach. Canopus leverages the two-qubit canonical representation and the monodromy polytope theory to model the synthesis cost for more intelligent $\mathrm{SWAP}$ insertion during qubit routing. We also formalize the commutation relations between two-qubit gates through the canonical form, providing a generalized approach to commutativity-based optimization. Experiments show that Canopus consistently reduces routing overhead by 15%-35% compared to state-of-the-art methods across various backend ISAs and device topologies. More broadly, this work establishes a coherent method for co-exploration of program patterns, quantum ISAs, and hardware topologies, yielding concrete guidelines for hardware-software co-design. This is the first practical demonstration of how to efficiently utilize advanced quantum ISAs, opening the door to more powerful and synergistic quantum systems.
@@ 1. Introduction
Holistic benchmarks of quantum computers such as quantum volume (cross2019validating) are predicated on concurrent advancements in both hardware and software. Recently, numerous systematic techniques regarding compiler optimization and architecture design have been presented to push the limit of hardware performance. Quantum compilers play a pivotal role in this process, translating high-level programs into executable single-qubit (1Q) and two-qubit (2Q) gates on realistic quantum hardware. This typically involves several stages: (1) compiling programs into basic quantum gates, (2) performing hardware-agnostic (logical-level) circuit optimization, (3) resolving backend topology constraints via qubit placement and routing, and (4) converting circuits to native gates for further optimization and scheduling. The primary goal of compiler optimization is to lower the 2Q gate count and circuit depth while resolving backend constraints, with a particular emphasis on 2Q gates due to their significantly higher error rates compared to 1Q gates.
@@
Most studies on qubit routing rely on a simplified routing model, where circuit cost is quantified by the $\mathrm{CX}$-based gate count and circuit depth while each $\mathrm{SWAP}$ gate is unrolled into three $\mathrm{CX}$ gates according to the textbook pattern $\mathrm{SWAP}_{q_{0},q_{1}}=\mathrm{CX}_{q_{0},q_{1}}\mathrm{CX}_{q_{1},q_{0}}\mathrm{CX}_{q_{0},q_{1}}$. However, this $\mathrm{CX}$-centric view is misaligned with the physical reality of modern quantum devices. Although quantum algorithms are typically expressed in terms of $\mathrm{CX}$ gates, the underlying hardware may not execute native $\mathrm{CX}$-equivalent gates, nor does this gate/circuit cost quantification method accurately reflect the true operational cost. Indeed, beyond the native support for $\mathrm{CX}$-equivalent gates (e.g., $\mathrm{CZ}$ (krantz2019quantum), Cross-Resonance (rigetti2010fully), Mølmer-Sørensen (bruzewicz2019trapped)), modern quantum hardwares increasingly feature diverse native 2Q basis gates in recent years. These alternative basis gates, or the abstracted instruction set architectures (ISAs) can be more powerful than $\mathrm{CX}$-equivalent gates in terms of synthesis capabilities and fidelity of realization, such as $\mathrm{\sqrt{iSWAP}}$ (huang2023quantum), the $\mathrm{iSWAP}$-family and $\mathrm{CX}$-family gates (mckinney2024mirage; qiskitXXDecomposer), and heterogeneous basis gates (peterson2022optimal; mckinney2024mirage). With such ISAs, $\mathrm{SWAP}$ can be implemented with a lower cost than three $\mathrm{CX}$ gates or even be natively realized with high fidelity (wei2024native; chen2025efficient; nguyen2024programmable). Therefore, the simplified routing model completely ignores the backend ISA properties, severely limiting the potential of compiler optimization. Furthermore, the absence of systematic compiler optimization methods across these diverse (even complex, heterogeneous) ISAs has prevented the community from fully exploiting their power and exploring the rich software-hardware co-design space.
@@
The main ideas of Canopus are as follows: ① Significant optimization opportunities arise from incorporating the cost of native-gate synthesis directly into the qubit routing process. For instance, synthesizing a 2Q block and a subsequent $\mathrm{SWAP}$ on the same qubit pair as a single composite operation is often more efficient than synthesizing them individually. ② Expanding the quantum ISA is crucial to boost the performance of real-world quantum applications. For example, the fractional $\mathrm{ZZ}(\theta)$ gate set widely adopted by hardware vendors (e.g., IBM (ibmFractionalGates), Quantinuum (quantinuumArbitraryAngleGates), IonQ (ionqPartialGates)) enables more efficient execution of chemistry simulation kernels within which many 2-local Pauli rotations are involved. The combination of $\mathrm{CX}$ and $\mathrm{iSWAP}$ gates have been demonstrated to benefit stabilizer circuits to protect error-corrected qubit information (zhou2024halma). ③ The monodromy polytope theory (peterson2020fixed) based on the canonical representation of 2Q gates (zhang2003geometric) provides a formal, universal, and quantitative description of the 2Q synthesis cost for arbitrary quantum ISAs. This formalism enables unified compiler optimization. With these insights in mind, Canopus performs intelligent $\mathrm{SWAP}$ insertion during qubit routing to holistically minimize post-mapping circuit cost (in terms of both gate count and depth) given any quantum ISA, thus performing deep routing-synthesis co-optimization and resulting in significantly lower routing overhead. Importantly, although Canopus is ISA-aware, it always operates on the canonical-form circuits, and the gate/circuit cost quantification via monodromy polytope is independent of the specific ISA rebase implementation. In this sense, Canopus offers LLVM-style compiler optimization.
Experimental results demonstrate that Canopus consistently provides 15%-35% reduction (in terms of both gate count and depth) of routing overhead compared to other SOTA methods across representative quantum ISAs, including the conventional CX ISA. This cross-ISA comparison also reveals some consistent or program-specific and topology-specific guidelines for hardware-software co-design. Source code and data are available via the Github link. Our work makes the following key contributions:
❶We utilize the canonical 2Q gate representation and the monodromy polytope to quantify costs of 2Q gates and the overall circuit. This formal approach accurately guides synthesis-routing co-optimization and cross-ISA evaluation.
@@
Our case studies, including the real-machine QFT kernel execution and the end-to-end QEC circuit simulation, unequivocally showcase Canopus’ superiority in both near-term and fault-tolerant applications. For example, on the task of mapping QFT on 1D chain topology, Canopus finds the provably optimal routing scheme, surpassing the results previously reported as optimal in prior work (zhang2021time); and experiments on IBM’s QPUs demonstrate that, compared to Qiskit, Canopus reduces errors by an average of 26.89% and 34.98% for the $\mathrm{CZ}$ and $\mathrm{ZZ}(\theta)$ gate sets, respectively.
@@ 2.2. Canonical description of 2Q gates
Figure 3. Geometric illustration of canonical gates confined to the Weyl chamber. For visualization convenience, herein the Weyl chamber is confined to $\left\{\frac{\pi}{4}\geq x\geq y\geq z\geq 0\right\}\cup\left\{\frac{\pi}{4}\geq\frac{\pi}{2}-x\geq y\geq z\geq 0\right\}$, equivalent to the canonical coefficient convention $\left\{(a,b,c)\,|\,\frac{1}{2}\geq a\geq b\geq|c|\right\}$.
@@
$\mathrm{Can}(a,b,c):=e^{-i\frac{\pi}{2}(a\,XX+b\,YY+c\,ZZ)},\,\frac{1}{2}\geq a\geq b\geq\lvert c\rvert$
sandwiched by local 1Q gates such that we say $U$ is locally equivalent to ($\sim$) the canonical form $\mathrm{Can}(a,b,c)$.
@@ 2.3. Gate realization cost on hardware
The transformed circuits via qubit routing will be ultimately converted into basis gates for execution on hardware. Basis gates refer to those natively implemented and calibrated on physical platforms. Typical native gates in superconducting are $\mathrm{CR}$ (rigetti2010fully), $\mathrm{CZ}$, and $\mathrm{iSWAP}$ gates (krantz2019quantum). The realization cost of basis gates involves multiple aspects, including the benchmarked fidelity, gate duration, calibration efficiency, etc. For example, gates with shorter duration are more likely to achieve high fidelity, as qubit decoherence dominates the noise source; although some gate schemes can now implement more basis gates (chen2025efficient; nguyen2024programmable), those with simpler pulse control are more likely to be calibrated with high precision, such as the $\mathrm{iSWAP}$-family gates on flux-tunable transmons.
2Q gates that are not natively implemented must be synthesized by native gates. Their realization cost is determined by the basis gates required for synthesis. For example, any 2Q gate can be minimally synthesized by 3 $\mathrm{CX}$ gates, except for $\mathrm{Can}(a,b,0)$ for which the required $\mathrm{CX}$ count is 2. Conventionally, $\mathrm{SWAP}$ is regarded as 3 times that of $\mathrm{CX}$ realization cost, while it can also be synthesized by “1 $\mathrm{CX}$ + 1 $\mathrm{iSWAP}$” or “3 $\mathrm{\sqrt{iSWAP}}$” gates. The monodromy polytope theory was proposed to determine the optimal synthesis cost for any 2Q gate given a specific set of basis gates through analysis of local invariants of canonical gates (peterson2020fixed). By this method, the set of gates realizable by a specified number of 2Q gates from the basis set, with arbitrary 1Q gates, corresponds to a polytope within the Weyl chamber. For instance, the polytope for 2 $\mathrm{\sqrt{iSWAP}}$ gates is a tetrahedron confined to $\left\{1/2\geq a\geq b+\lvert c\rvert\right\}$ (huang2023quantum).
@@ 3. Motivation
Limitations of conventional qubit routing models
@@
Co-optimization as the key to unlocking the superiority of advanced ISAs
@@
Prior efforts have been narrowly focused on local 2Q or multi-qubit synthesis tasks (huang2023quantum; tang2024quantum) or brute-force numerical optimizations (davis2019heuristics; younis2021qfast). Such rebase passes are tailored to a specific quantum ISA and fail to provide clear benefits when applied to advanced ISAs in realistic workloads (kalloor2024quantum). This has led to a critical question lingering in the community: “Are these more expressive, noise-resilient ISAs actually better?” Recently there have been attempts to harness the properties of advanced ISAs, although through manual and unsophisticated design, such as the $\mathrm{\sqrt{iSWAP}}$-based routing-synthesis optimization (mckinney2024mirage) and the $\mathrm{CX}$-$\mathrm{iSWAP}$ based routing for defect effect mitigation (zhou2024halma). In our work, we highlight that collaborative compiler optimization, especially at the stage following logical-level circuit optimization and followed by the final ISA rebase pass, is a key to fully exploit capabilities of those powerful ISAs: First, high-level algorithms are expressed in the $\mathrm{CX}$ representation, which then undergo template-based and peephole optimizations that are highly sophisticated and tailored for $\mathrm{CX}$-based circuit patterns (e.g., commutativity, Clifford equivalence); Second, the gap between the naïve qubit routing model and backend ISA properties apparently leaves a large untapped co-optimization space for both qubit routing and ISA rebase. Thus we are committed to validating this point through a systematic ISA-aware routing framework.
The “Tower of Babel dilemma” for utilizing diverse ISAs
The proliferation of diverse quantum ISAs—from monolithic to complex, heterogeneous basis gate sets supported by various physical platforms—has created a “Tower of Babel dilemma” in the architecture and systems community. Developing bespoke compiler optimizations for each unique hardware backend is unsustainable, leading to the same software fragmentation that we have encountered in classical computing. Consequently, it is important to seek a unified approach that can effectively handle various platform-specific abstractions resembling the LLVM compiler (Lattner2004). The recently proposed monodromy polytope theory (peterson2020fixed), for example, provides the the method for optimal analysis of ISA synthesis capabilities. Specifically regarding the circuit-level compiler optimization, the monodromy polytope with canonical 2Q gate representations offers a unified approach to evaluating circuit cost and modeling routing-synthesis co-optimization. Building on this, Canopus proves to be an elegant and unified solution to the Tower of Babel dilemma at the compiler level.
Coherent cross-ISA, topology, and program pattern co-exploration
@@ 4.1. Overview
The overall routing procedure of Canopus is illustrated in Figure 4. The input circuit is rebased to $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ before being fed to the routing pass. All processes operate on the directed acyclic graph (DAG) representation of the circuit. Canopus integrates the ISA-specific synthesis cost model into its $\mathrm{SWAP}$ search process and determines the most appropriate $\mathrm{SWAP}$ at each route step. The routing cost is efficiently computed by a formal analysis of 2Q canonical forms, without the ISA rebase step. Thus, the output is still a circuit DAG represented in $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ with inserted $\mathrm{SWAP}$ gates.
Notably, Canopus inherits the basic concepts and data structures introduced in Sabre (li2019tackling), which is one of the industrial-standard qubit mapping algorithms. Given the input circuit DAG, Canopus first attempts to map 2Q gates layer by layer via extracting the front layer $F$, peeling executable gates and searching $\mathrm{SWAP}$ gates to minimize the routing cost according to a heuristic cost function. On the backbone of Sabre, we further introduce several key data structures such as the last mapped layer $L$ and the wire duration record $D$, to support the efficient implementation of ISA-aware routing and reducing both gate count and depth related routing overhead.
@@
As introduced in Section 2.3, given any basis gate set, the synthesis cost of a target 2Q gate can be exactly computed through monodromy polytope (peterson2020fixed). This cost (which basis gates are sufficient for synthesis) only depends on the canonical coefficients of the target gate. For example, Figure 5 illustrates various polytopes for the gate set $\bigl\{\mathrm{\sqrt{iSWAP}},\,\mathrm{ECP}\bigr\}$. In practice, the costs of each basis gate are pre-defined, thus the whole set of polytopes helps decide the optimal synthesis scheme with the minimal circuit cost we should prioritize. For example, if $\mathrm{\sqrt{iSWAP}}$ and $\mathrm{ECP}$ have the same unit cost, the $\mathrm{SWAP}\sim\mathrm{Can}\bigl(\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)$ gate realization will prioritize the “1 $\mathrm{\sqrt{iSWAP}}$ + 1 $\mathrm{ECP}$” combination; if $\mathrm{ECP}$ cost is set to be more than twice that of $\mathrm{\sqrt{iSWAP}}$, the $\mathrm{SWAP}$ realization prioritizes the “3 $\mathrm{\sqrt{iSWAP}}$” pattern.